Spring is compressed to a length of 20 cm by a 20 kg object on a 20 degrees inclined surface with coefficient of kinetic friction 0.30. What is the initial velocity of the object moving up the plane? The spring constant is 980 N/m


1
Expert's answer
2021-10-19T15:08:11-0400

To release the object, the spring must overcome friction, perform work against gravity to lift the mass, and increase its potential energy:


"\frac12kx^2=\frac12mv^2+\mu mgx\cos\theta+mgx\sin\theta,\\\space\\
v=\sqrt{x\bigg[\frac{kx}{m}-2g(\mu cos\theta+\sin\theta)\bigg]},"

where

"x\bigg[\frac{kx}{m}-2g(\mu cos\theta+\sin\theta)\bigg]<0."

Therefore, the object won't move up the slope.